A & B take equal amount of loan. A pays 8% Interest two times per year while b pays 12% Interest three times per year. After 17 years, B pays 1400 more than that of A. What amount they take the loan at starting?
Answer by Boreal(15235)

A=Ao{1+r/n}^nt
{1+.08/2}^34+1200={1+.12/3}}^51
1.04^34+ 1200= {1/04}^51
3.7943 +1200=7.3909
In other words,
A for A=3.7943*Ao
A for B=7.3909*Ao
3.7943Ao+1400=7.3909Ao, since I need to add $1400 to A to match B.
1400=3.5967Ao
Ao=$389.25
389.25{1.04}^34=$1476.94
389.25(1.04)^51=2876.93, $1400 more within rounding error.
The answer is $389.25